What is meant by bending moment? Calculate the bending moment of a thin uniform bar of rectangular and circular cross-sections.  


1
Expert's answer
2021-03-04T11:51:23-0500

A bending moment is the reaction induced in a structural element when an external force or moment is applied to the element, causing the element to bend.

Bending moment of a thin uniform bar of rectangular and circular cross-section

Bending moment of a thin uniform bar of rectangular cross-section:


"M=\frac{bd^3}{12R}"

Bending moment of a thin uniform bar of circular cross-section:


"M=\frac{\pi Yr^4}{4R}"